### v2

# Evaluating Sparse Autoencoders for Monosemantic Representation

Abstract
A key barrier to interpreting large language models is polysemanticity, where neurons activate for multiple unrelated concepts. Sparse autoencoders (SAEs) have been proposed to mitigate this issue by transforming dense activations into sparse, more interpretable features. While prior work suggests that SAEs promote monosemanticity, no quantitative comparison has examined how concept activation distributions differ between SAEs and their base models. This paper provides the first systematic evaluation of SAEs against base models through activation distribution lens. We introduce a fine-grained concept separability score based on the Jensen–Shannon distance, which captures how distinctly a neuron’s activation distributions vary across concepts. Using two large language models (Gemma-2-2B and DeepSeek-R1) and multiple SAE variants across five datasets (including word-level and sentence-level), we show that SAEs reduce polysemanticity and achieve higher concept separability. To assess practical utility, we evaluate concept-level interventions using two strategies: full neuron masking and partial suppression. We find that, compared to base models, SAEs enable more precise concept-level control when using partial suppression. Building on this, we propose Attenuation via Posterior Probabilities (APP), a new intervention method that uses concept-conditioned activation distributions for targeted suppression. APP achieves the smallest perplexity increase while remaining highly effective at concept removal  [^1] Code will be released publicly after paper acceptance..
Evaluating Sparse Autoencoders for Monosemantic Representation
Moghis Fereidouni, Muhammad Umair Haider, Peizhong Ju, A.B. Siddique, University of Kentucky, moghis.fereidouni@uky.edu, muhammadumairhaider@uky.edu, peizhong.ju@uky.edu, siddique@cs.uky.edu
1 Introduction
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, often matching or surpassing human-level performance Luo et al. (2025); Achiam et al. (2023); Touvron et al. (2023); Guo et al. (2025); Eslamian and Cheng (2025). Nonetheless, understanding how these models internally represent and manipulate concepts remains a major challenge. A key obstacle is polysemanticity; the phenomenon where individual neurons respond to multiple, semantically distinct concepts rather than encoding single, interpretable features Janiak et al. (2023); Olah et al. (2017); Nguyen et al. (2016). This entanglement complicates the interpretation and analysis of model behavior, posing a significant barrier to building transparent and controllable AI systems Sharkey et al. (2025); Marshall and Kirchner (2024); Bereska and Gavves (2024).
Dictionary learning via sparse autoencoders (SAEs) Huben et al. (2024); Gao et al. (2025) has recently emerged as a promising approach to mitigating polysemanticity in neural representations. SAEs aim to transform dense activations of a desired component of the base LLM into sparse features by enforcing sparsity and encouraging each neuron to specialize in distinct, concept-specific features Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). The goal is to produce monosemantic representations, where individual neurons respond to single, well-defined concepts Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). The underlying hypothesis is intuitive; if we can force the model to use fewer neurons simultaneously, each active neuron should correspond to a more distinct and interpretable concept. Empirical studies have shown that SAEs can uncover interpretable features across domains such as vision and language, facilitating improved interpretability Shu et al. (2025); Huben et al. (2024); Pach et al. (2025).
Most existing evaluations of SAE interpretability are qualitative, relying on case studies or anecdotal neuron visualizations that provide limited systematic insight Kissane et al. (2024); Li et al. (2025). The remaining quantitative efforts mainly examine whether neurons are active for given concepts, rather than capturing the full distributional structure of concept activations across neurons Minegishi et al. (2025); Karvonen et al. (2025).
In this work, we conduct a systematic investigation into the effectiveness of SAEs in promoting monosemanticity in the internal representations of LLMs through the distributional lens. Particularly, we conduct comprehensive evaluations using two large language models (Gemma-2-2B Rivière et al. (2024) and DeepSeek-R1 Guo et al. (2025)) and various SAEs of different widths and sparsity levels on five benchmark datasets, including both word-level (POS tagging and NER) and sentence-level (e.g., AG News) tasks. We begin by quantifying polysemanticity using overlap statistics, measuring the fraction of salient neurons that respond to multiple, semantically distinct concepts. While SAEs exhibit lower polysemanticity than their base models, this overlap-based analysis remains coarse-grained; it treats all neurons that respond to multiple concepts as equally entangled, without considering how their activations vary across those concepts. In practice, a neuron may activate for several concepts, yet do so with clearly distinct activation distributions, suggesting behavior that may still be considered monosemantic. That is, monosemanticity is not solely about binary activation overlap, but rather about the separability of a neuron’s activation distributions across concepts.
We formalize this view by introducing a new concept separability score, based on the Jensen–Shannon distance Lin (1991). This fine-grained, distribution-aware metric quantifies how well a neuron’s activations separate across different concepts by measuring the distance between their activation distributions. Using this score, we find that SAEs exhibit higher concept separability than their dense counterparts.
To further evaluate the practical utility of monosemantic representations of SAEs, we examine their effectiveness in enabling concept-level model interventions. Specifically, we assess how precisely concept-related behavior can be suppressed in SAEs compared to base models. We evaluate two intervention strategies; full neuron masking, which suppresses all activations of salient neurons associated with a target concept, and partial suppression, which intervenes in activations selectively based on their distributional association with the concept. Across SAEs and the base model, partial suppression outperforms full masking in most cases, achieving more effective suppression of the target concept while better preserving unrelated model behavior. Furthermore, our results show that SAEs support more precise and effective concept removal than their dense counterparts, especially when applying partial suppression methods. These findings reinforce the idea that concept separability, when defined in terms of activation distributions rather than binary activation overlap, offers better model control and interpretability.
Additionally, motivated by the varying separability of concept activations across neurons, we introduce Attenuation via Posterior Probabilities (APP), a new intervention method that leverages concept-conditioned activation distributions to selectively suppress target concepts with minimal side effects. Specifically, APP computes the posterior probability that a given activation corresponds to a target concept and attenuates it accordingly. Among all methods evaluated, APP achieves the smallest degradation in language modeling quality (lowest perplexity increase) while remaining highly competitive with other baselines in targeted concept removal across both SAEs and base models.
In summary, this work makes the following contributions:
  • We present the first quantitative analysis of monosemantic representations in SAEs relative to their base LLM through distributional lens.
  • We introduce a concept separability score, based on the Jensen-Shannon distance, a fine-grained, distribution-aware metric that captures how well neuron activations separate for different concepts.
  • We propose a new intervention method, which is the least invasive concept erasure technique and highly competitive with existing methods in removing the targeted concept.
2 Preliminaries
Neuron. A neuron refers to a component of a hidden state vector in a transformer layer. Given a hidden state $h^{l}\in\mathbb{R}^{d}$ at layer $l$, the $j$-th neuron is denoted by $h^{l}_{j}$.
Concept. A concept $c_{i}\in C$ is a semantic category assigned to each input (or its components), where $C={c_{1},\dots,c_{k}}$. For instance, sentence-level types (e.g., declarative, interrogative) or word-level tags (e.g., noun, verb) can serve as concepts. In this work, we focus both on sentence-level and word-level concepts.
Datasets and Models. We use five datasets: Part-of-Speech Tagging (POS) Pasini et al. (2021), Named Entity Recognition (NER) Ding et al. (2021), AG News Zhang et al. (2015), Emotions Saravia et al. (2018), and DBpedia Zhang et al. (2015). Our analysis is conducted using two large language models: (1) the Gemma-2-2B model Rivière et al. (2024) along with its corresponding Sparse Autoencoders (SAEs) from GemmaScope Lieberum et al. (2024), and (2) the DeepSeek-R1 model Guo et al. (2025) with its associated SAE from LlamaScope He et al. (2024).
Figure 1: SAEs reduce neuron overlap in comparison to the base model, indicating lower polysemanticity. Higher-capacity SAEs (65k) further reduce overlap, suggesting more effective assignment of distinct neurons to separate concepts.
3 Analyzing Polysemanticity in SAEs
A neuron is considered polysemantic when it responds to multiple, distinct concepts rather than a single, well-defined one. Several studies have demonstrated that such polysemantic behavior is common in neural networks Elhage et al. (2022); Bau et al. (2017); Scherlis et al. (2022); Lecomte et al. (2024); Marshall and Kirchner (2024); Olah et al. (2017); Nguyen et al. (2016). This polysemanticity reduces the interpretability of models, motivating the development of Sparse Autoencoders (SAEs) Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). SAEs are designed to encourage sparsity in neural activations, aiming to align each neuron with a specific, distinct concept and thereby promote monosemanticity and interpretability. In the following, we evaluate SAEs to better understand their effectiveness in improving monosemanticity.
Figure 2: Across base model and SAEs (SAE-16k, SAE-65k), neurons exhibit varying degrees of separability in their activations. Some have completely overlapping activations across concepts, others show partial or clear separation. This variability underscores the importance of using distribution-aware metrics when assessing monosemanticity.
3.1 Salient Neuron Overlap
As a first step, we quantify polysemanticity by measuring the overlap percentage of salient neurons (i.e., neurons with high mean activation) across concepts. Specifically, for each concept, we identify the top 80 salient neurons by mean activation. Then, the overlap percentage is computed as the intersection-over-union of these top-K sets across concepts. This metric captures the extent to which neurons are shared across concepts, reflecting shared saliency and polysemanticity. The Figure 1 compares this shared saliency for the base model and two Sparse Autoencoders (SAEs) with different latent dimensions (16k and 65k), while maintaining comparable sparsity levels (116 vs. 93 active neurons). In the DBpedia dataset, for instance, we observe that nearly 9% of the top-activated (salient) neurons in the base model are shared across all 14 concepts. Moreover, the Figure 1 confirms that SAEs exhibit reduced conceptual overlap, suggesting less polysemanticity; however, polysemantic neurons are still present. For example, in the POS dataset, the percentage of shared salient neurons drops from around 38% in the base model to approximately 18% in the SAE with 16k dimensions. This is still a relatively high percentage, indicating that over 18% of the most salient neurons are shared across all four concepts. Moreover, another notable observation is that the 65k-dimensional SAEs exhibit lower polysemanticity than their 16k-dimensional counterparts across all datasets. This reinforces the idea that larger SAEs have greater capacity to allocate distinct neurons to specific concepts, thereby enhancing interpretability. For additional analysis evaluating all active SAE neurons, not just the top 80, see Appendix E.
3.2 SAEs Activation Distributions
So far, we have analyzed polysemanticity by quantifying neuron overlap based on activation frequency, specifically, how often a neuron is active or salient across different concepts. While these approaches offer useful aggregate insights, they treat neuron activation as a binary or averaged signal, overlooking the distributional characteristics of how neurons respond to each concept. In other words, a neuron might be shared across multiple concepts, but the manner in which it activates for each could vary significantly, ranging from broad, overlapping responses to distinct, well-separated patterns. To better capture this nuance, Figure 2 displays the full activation distributions of selected neurons across concept classes in the AG News dataset, comparing the base model with two SAE variants of differing capacities. We include an analogous visualization for the NER dataset in Appendix B.
These plots reveal two key patterns: (1) consistent with prior observations Haider et al. (2025), neuron activations in both the base model and SAEs tend to follow approximately Gaussian distributions; and (2) while some neurons are shared across multiple concepts, their activation distributions can range from highly overlapping to clearly separable. This underscores a fundamental limitation of mean-based or binary overlap measurements, which can overlook meaningful distinctions in activation behavior. To more accurately measure polysemanticity, it is therefore necessary to employ a metric that captures the full shape of activation distributions across concepts. To this end, we introduce a new concept separability score, based on the Jensen-Shannon distance, which quantifies the degree of separation between concept-specific activation distributions.
POS AG News Emotions DBpedia NER
Base 0.343 0.366 0.322 0.405 0.308
SAE
(width_16k/l0_116)
0.539 0.650 0.431 0.621 0.581
SAE
(width_65k/l0_93)
0.600 0.709 0.446 0.680 0.621
Table 1: Separability Score $S$ across datasets.
Concept Separability Score. To quantify how separable a neuron’s activation distributions are across $k$ concepts, we first define the probability density function as:
$f_{h^{l}_{j}\mid c_{i}}(x)=p\bigl(h^{l}_{j}=x\mid c_{i}\bigr),\quad i=1,\dots,k,$
with this density function at hand, we define mixture as $M^{(l,j)}(x)=\frac{1}{k}\sum_{i=1}^{k}f_{h^{l}_{j}\mid c_{i}}(x)$,
and then compute the generalized Jensen–Shannon divergence Lin (1991) as
$\mathrm{JSD}\bigl(f_{h^{l}_{j}\mid c_{1}},\dots,f_{h^{l}_{j}\mid c_{k}}\bigr)=H\bigl(M^{(l,j)}\bigr)-\frac{1}{k}\sum_{i=1}^{k}H\bigl(f_{h^{l}_{j}\mid c_{i}}\bigr).$
Next, we take the square root of the JSD and normalize it to obtain a proper distance metric bounded within $[0,1]$:
$D_{\mathrm{JS}}\bigl(f_{h^{l}_{j}\mid c_{1}},\dots,f_{h^{l}_{j}\mid c_{k}}\bigr)=\frac{\sqrt{\mathrm{JSD}\bigl(f_{h^{l}_{j}\mid c_{1}},\dots,f_{h^{l}_{j}\mid c_{k}}\bigr)}}{\sqrt{\log_{2}k}},$
Moreover, we assign $D_{\mathrm{JS}}=1$ whenever a neuron’s activations are all attributed to a single concept. Finally, the layer‐level separability score is
$S^{l}=\frac{1}{d}\sum_{j=1}^{d}D_{\mathrm{JS}}\bigl(f_{h^{l}_{j}\mid c_{1}},\dots,f_{h^{l}_{j}\mid c_{k}}\bigr),$
where $d$ is the number of neurons in layer $l$.
Table 1 reports the concept separability score $S\in[0,1]$ for Gemma, where higher values indicate more distinct activation distributions across concepts, indicating higher monosemanticity (Full JS scores for all settings and both LLMs are in Appendix G). Across all five datasets, SAEs substantially improve separability over the base model. For instance, on DBpedia (14 classes), the score increases from 0.405 in the base model to 0.621 and 0.680 for the 16K- and 65K-dimensional SAEs, respectively, an increase of over 50%. Furthermore, higher-capacity SAEs consistently yield greater separability, supporting the view that more expressive latent spaces better disentangle concepts.
Building on our finding that sparse autoencoders yield more separable concept representations, we next assess whether this improved separability enables more precise concept erasure interventions.
4 Concept Erasure
Concept erasure encompasses interventions that aim to remove a specific concept from a model’s internal representation, ideally without affecting other concepts Dalvi et al. (2019a, b); Dai et al. (2022); Morcos et al. (2018). Formally, let $M$ be a trained model that maps an input $x$ to a concept label $M(x)=c$. An ideal erasure yields a modified model $M^{\prime}_{\mathrm{ideal}}$ satisfying:
$M^{\prime}_{\mathrm{ideal}}(x)=\begin{cases}\neq M(x),&\text{if }M(x)=c,\\ =M(x),&\text{if }M(x)\neq c.\end{cases}$
That is, the model should unlearn the target concept $c$ while preserving its behavior on all other concepts.
As shown in Figure 2, some neurons exhibit considerable overlap in their activation distributions across concepts, while others show separability. This pattern appears in both the base model and SAEs, suggesting that concept erasure techniques should not treat all activation values in one neuron identically. Instead, we propose a more targeted approach that considers where an activation falls within the distribution. Specifically, values in regions uniquely tied to a concept (i.e., those regions of distribution that are clearly separable from other concepts) should be suppressed more strongly, while regions shared across concepts should be dampened more conservatively to preserve other concepts. To achieve this, we introduce Attenuation via Posterior Probabilities (APP), which modulates suppression based on distributional separability.
4.1 Attenuation via Posterior Probabilities (APP)
Given all neurons $h^{l}_{j}$ of layer $l$, with individual activations $x^{l}_{j}$ for $j=1,\dots,d$, and a target concept $c_{i}\in C$, our goal is to selectively suppress the activation $x^{l}_{j}$ that is attributable to $c_{i}$, while preserving contributions from other concepts. We begin by computing the posterior probability that a given activation $x^{l}_{j}$ arose from concept $c_{i}$, under the assumption that all concepts are a priori equally likely:
$\begin{aligned} \pi_{j,i}(x^{l}_{j})&=p\bigl(c_{i}\mid h^{l}_{j}=x^{l}_{j}\bigr)=\frac{p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr)\,p(c_{i})}{\displaystyle\sum_{m=1}^{k}p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr)\,p(c_{m})},\\ &=\frac{p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{i}\bigr)}{\displaystyle\sum_{m=1}^{k}p\bigl(h^{l}_{j}=x^{l}_{j}\mid c_{m}\bigr)}\;\equiv\;\frac{f_{h^{l}_{j}\mid c_{i}}(x^{l}_{j})}{\displaystyle\sum_{m=1}^{k}f_{h^{l}_{j}\mid c_{m}}(x^{l}_{j})}.\end{aligned}$
By definition, $\sum_{i=1}^{k}\pi_{j,i}(x^{l}_{j})=1$.
To avoid unreliable posterior estimates from low-density regions, we limit our attention to the central region of the target concept’s activation distribution, where density estimates are more reliable. Let $\mu_{j,i}$ and $\sigma_{j,i}$ denote the mean and standard deviation of neuron $h^{l}_{j}$ under concept $c_{i}$, and define the valid damping window as:
$W_{j,i}\;=\;\bigl[\mu_{j,i}-2.5\,\sigma_{j,i},\;\mu_{j,i}+2.5\,\sigma_{j,i}\bigr].$
The damping factor $\alpha_{j,i}(x)$ is then defined as:
$\alpha_{j,i}(x)=\begin{cases}1-\pi_{j,i}(x),&x\in W_{j,i},\\ 1,&\text{otherwise}.\end{cases}$
With $\alpha\approx 0$ when $x$ is very typical of the target concept $c_{i}$. Finally, we apply this factor to dampen the activation:
$\widetilde{x}^{l}_{j}=\alpha_{j,i}(x^{l}_{j})\,x^{l}_{j}=\begin{cases}\bigl[1-\pi_{j,i}(x^{l}_{j})\bigr]\,x^{l}_{j},&|x^{l}_{j}-\mu_{j,i}|\leq 2.5\,\sigma_{j,i},\\[3.0pt] x^{l}_{j},&\text{otherwise}.\end{cases}$
This formulation enables precise, concept-aware suppression while leaving unrelated or uncertain activations unchanged.
4.2 Baseline Methods
To comprehensively evaluate concept erasure effectiveness, we compare APP (which is a partial suppression method) against three other partial methods and one full-masking baseline.
AURA Suau et al. (2024): Ranks neurons by AUROC, selects those with $\mathrm{AUROC}>0.5$, and dampens their output based on $\mathrm{AUROC}$.
Range Masking Haider et al. (2025): The activations of concept-relevant neurons (highly activated) are suppressed when they fall within their typical range ($\mu\pm 2.5\sigma$).
Adaptive Dampening Haider et al. (2025): The activations of concept-relevant neurons (highly activated) are dampened in proportion to their distance from the concept mean.
Full Masking Dalvi et al. (2019a); Dai et al. (2022); Antverg and Belinkov (2022): concept-relevant neurons (highly activated) are fully zeroed out to eliminate the target concept.
Table 2: Concept Erasure Results by Method and Model Type across Datasets (Gemma-2-2B). Bolded values indicate the best performance, and underlined values denote the second-best. Results are grouped by intervention method (e.g., APP, AURA, Adaptive) and model type (Base vs. SAE variants) across five benchmark datasets.
Type Method POS AG News Emotions DBpedia NER
$\Delta_{\mathrm{Acc}}\uparrow$ $\Delta_{\mathrm{Conf}}\uparrow$ $\mathrm{DPPL}\downarrow$ $\Delta_{\mathrm{Acc}}\uparrow$ $\Delta_{\mathrm{Conf}}\uparrow$ $\mathrm{DPPL}\downarrow$ $\Delta_{\mathrm{Acc}}\uparrow$ $\Delta_{\mathrm{Conf}}\uparrow$ $\mathrm{DPPL}\downarrow$ $\Delta_{\mathrm{Acc}}\uparrow$ $\Delta_{\mathrm{Conf}}\uparrow$ $\mathrm{DPPL}\downarrow$ $\Delta_{\mathrm{Acc}}\uparrow$ $\Delta_{\mathrm{Conf}}\uparrow$ $\mathrm{DPPL}\downarrow$
Base APP 0.224 0.210 0.571 0.347 0.266 0.394 0.051 0.095 0.151 0.113 0.104 0.076 0.079 0.129 0.201
Aura 0.063 0.091 1.195 0.112 0.092 2.130 -0.019 0.004 2.273 0.092 0.064 2.792 0.033 0.044 1.515
Range 0.151 0.1006 2.258 0.081 0.098 1.026 0.036 0.104 0.784 0.168 0.109 0.700 0.135 0.209 1.978
Adaptive 0.157 0.129 1.158 0.050 0.089 0.545 0.032 0.090 0.428 0.120 0.099 0.387 0.095 0.173 1.052
Full 0.115 0.069 22.873 0.096 0.080 16.914 0.038 0.113 17.438 0.169 0.102 10.922 0.134 0.176 23.598
SAE width: 65k l0: 93 APP 0.302 0.270 0.342 0.601 0.442 0.230 0.231 0.279 0.075 0.399 0.219 0.114 0.507 0.520 0.161
Aura 0.406 0.312 0.360 0.577 0.367 0.592 0.336 0.311 0.286 0.357 0.213 0.490 0.519 0.514 0.359
Range 0.052 0.055 0.741 0.587 0.125 0.542 0.146 0.024 0.413 0.252 0.056 0.446 0.497 0.549 0.433
Adaptive 0.095 0.105 0.533 0.592 0.215 0.358 0.182 0.084 0.226 0.222 0.083 0.279 0.494 0.557 0.308
Full 0 0.0003 1.712 0.602 0.006 4.763 0.120 0.002 5.462 0.275 0.018 3.892 0.382 0.391 3.046
SAE width: 16k l0: 116 APP 0.216 0.191 0.468 0.582 0.369 0.569 0.265 0.274 0.166 0.312 0.152 0.140 0.377 0.397 0.531
Aura 0.250 0.192 0.577 0.596 0.286 1.262 0.339 0.280 0.493 0.257 0.140 0.836 0.389 0.415 0.741
Range 0.061 0.046 0.925 0.505 0.103 0.879 0.099 0.009 0.418 0.208 0.053 0.446 0.298 0.305 0.915
Adaptive 0.097 0.085 0.655 0.508 0.163 0.676 0.128 0.047 0.267 0.177 0.071 0.288 0.329 0.347 0.721
Full 0 0.0005 3.738 0.462 0.001 5.882 0.048 0.0005 6.454 0.187 0.0007 5.639 0.262 0.256 3.872
SAE width: 65k l0: 197 APP 0.350 0.322 0.398 0.615 0.458 0.433 0.247 0.285 0.406 0.362 0.235 0.083 0.530 0.654 0.346
Aura 0.384 0.317 0.576 0.611 0.397 0.940 0.311 0.327 0.584 0.297 0.235 0.596 0.567 0.681 0.510
Range 0.046 0.059 0.902 0.346 0.060 0.748 0.135 0.024 0.729 0.292 0.106 0.335 0.503 0.662 0.613
Adaptive 0.110 0.130 0.634 0.345 0.130 0.554 0.170 0.110 0.574 0.257 0.147 0.205 0.503 0.656 0.815
Full 0 0.00001 5.599 0.344 0.0007 7.422 0.110 0.002 7.613 0.216 0.019 6.203 0.292 0.384 3.625
SAE width: 16k l0: 285 APP 0.382 0.306 0.634 0.628 0.356 0.882 0.212 0.228 0.492 0.274 0.147 0.167 0.423 0.482 0.370
Aura 0.433 0.359 0.778 0.549 0.346 1.671 0.286 0.300 0.788 0.158 0.111 0.975 0.428 0.469 0.718
Range 0.051 0.047 1.356 0.456 0.069 1.852 0.111 0.012 1.395 0.181 0.057 0.958 0.338 0.392 0.924
Adaptive 0.179 0.148 0.975 0.441 0.149 1.312 0.121 0.072 0.948 0.145 0.083 0.582 0.376 0.435 0.669
Full 0 0.001 5.880 0.213 0.0004 16.284 0.062 0.001 8.117 0.153 0.023 9.510 0.167 0.191 8.661
4.3 Metrics
We evaluate the causal effect of our interventions using three metrics: task accuracy, predictive confidence, and perplexity.
Accuracy and confidence are measured both before and after intervention, for the target concept $c$ and all auxiliary concepts $c^{\prime}\neq c$. The goal is to assess how much the intervention selectively affects the target concept while minimizing disruption to others.
Let $D_{\mathrm{Acc}}$ denote the drop in accuracy for the target concept, and $D^{\prime}_{\mathrm{Acc}}$ the average drop in accuracy across auxiliary concepts. Similarly, let $D_{\mathrm{Conf}}$ and $D^{\prime}_{\mathrm{Conf}}$ be the drops in predictive confidence for the target and auxiliary concepts, respectively.
Using these, we compute two scores:
$\Delta_{\mathrm{Acc}}=D_{\mathrm{Acc}}-D^{\prime}_{\mathrm{Acc}},\quad\Delta_{\mathrm{Conf}}=D_{\mathrm{Conf}}-D^{\prime}_{\mathrm{Conf}}.$
Higher values of $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ indicate more precise interventions, strongly affecting the target concept while preserving performance on others. In the main text, we report only $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$; the full metric breakdowns are included in Appendix C.
Lastly, to capture the overall impact on the model’s generative ability, we measure the increase in perplexity:
$\mathrm{DPPL}=\mathrm{PPL}_{\mathrm{post}}-\mathrm{PPL}_{\mathrm{base}}.$
Comprehensive implementation details are presented in Appendix D, including our histogram-based KDE for concept-conditioned densities.
4.4 Results and Analysis
Comparison of Intervention Effectiveness: SAEs vs. Base Model. As it can be seen in Table 2, across nearly all settings, we find that partial intervention methods (particularly APP and AURA) consistently achieve higher $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ when applied to SAE representations compared to the base model. Specifically, for APP, SAE-based interventions outperformed the base model in 38 out of 40 comparisons. For AURA, SAE-based interventions were more effective in all 40 cases. In contrast, full masking shows less benefit from SAE representations; 19 out of 40 interventions resulted in better outcomes than when applied to the base model. This discrepancy suggests that coarse suppression methods fail to capitalize on the increased concept separability offered by SAEs. These findings reinforce that SAE representations are more disentangled and that fine-grained, distribution-aware methods are better equipped to exploit this structure for effective concept removal. Furthermore, focusing specifically on the APP intervention (which is a partial intervention), we observe that within the SAE family, increasing capacity consistently enhances intervention quality. In particular, $\Delta_{\mathrm{Conf}}$ consistently increases as we scale from 16 k to 65 k latent dimensions, reflecting improved confidence suppression for the target concept. $\Delta_{\mathrm{Acc}}$ also improves across most datasets, with only minor exceptions, further underscoring the role of latent dimensionality in enabling more precise and effective concept removal.
Partial Interventions Vs Full Interventions. Full masking ranks as the worst‐performing method in 36 out of 40 $\Delta$-metrics on the SAEs and it also produces the largest perplexity increase in all DPPL evaluations across SAEs and the base model. This underscores that distribution-aware partial methods (e.g., APP, AURA), which leverage activation distributions, are far more effective for targeted concept removal than the coarse, distribution-agnostic full-masking approach.
APP is least disruptive and highly competitive on concept erasure. APP consistently achieves the smallest $\mathrm{DPPL}$ across all 25 experiments, making it the least disruptive method. Adaptive follows as the second-best approach, attaining the second-smallest $\mathrm{DPPL}$ in 19 out of 25 cases. For concept removal, APP demonstrates strong and consistent performance; it most frequently achieves the best $\Delta_{Conf}$ (14/25 cases), while ranking first or second in 21 out of 25 settings. Regarding $\Delta_{Acc}$, AURA leads with 14 wins, though APP follows closely and ranks in the top two positions for 23 out of 25 settings. Overall, APP effectively removes target concepts while preserving predictive fluency better than all other baselines.
Comparative Analysis of APP and AURA. The superior effectiveness of AURA and APP on both $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ stems from the fact that they explicitly model not only the target‐concept distribution but also the distributions of all auxiliary concepts. By calibrating their interventions to maximize disruption of $c$ while minimizing collateral effects on $c^{\prime}\neq c$, both methods achieve higher $\Delta$ values than approaches that consider only the target distribution. Between these two, APP pulls ahead of AURA in terms of perplexity ($\mathrm{DPPL}$) because it leverages fine‐grained, activation‐specific damping rather than a single, per‐neuron factor. AURA mutes an “expert” neuron uniformly, regardless of whether a particular activation is highly characteristic of the target concept, whereas APP computes $\pi_{j,i}(x)$ for each activation and suppresses only the portions of the distribution uniquely associated with $c_{i}$. This activation-aware attenuation not only removes the targeted concepts effectively but also best preserves the model’s overall fluency, as evidenced by the smallest $\mathrm{DPPL}$.
Cross-Model Validation (DeepSeek). To verify that our observations are not specific to Gemma-2, we replicate experiments on DeepSeek-R1 (Appendix A). The results show the same trends, SAEs enable more selective concept removal than the base model. Moreover, APP achieves the smallest perplexity degradation ($\mathrm{DPPL}$) across all datasets except one, while remaining highly competitive on the concept-removal metrics ($\Delta_{\mathrm{Acc}}$, $\Delta_{\mathrm{Conf}}$).
4.5 Relation Between Erasure Methods and JS Distance
Figure 3: Separability Score vs. Erasure Ability (Partial)
Figure 4: Separability Score vs. Erasure Ability (Full)
Our distribution-aware separability score $S$ (see Subsection 1) serves as a natural predictor for the precision of concept erasure methods. Intuitively, the more separable a neuron’s activation distributions are, the easier it should be to suppress only the target concept while preserving unrelated behavior. We empirically validate this hypothesis in Figure 3, which plots our JS separability score $S$ (x-axis) against the change in accuracy difference $\Delta\text{Acc}$ (y-axis), aggregating data from both the Gemma and DeepSeek experiments (Both SAE and base models). A strong positive Pearson correlation ($r=0.771$, p < 0.001) emerges for the average performance of partial erasure methods (APP, AURA, Range, and Adaptive), confirming that higher separability reliably predicts more selective accuracy drops on the target concept. Detailed per-method correlations are provided in Appendix F. For comparison, we conducted the same analysis for the full masking approach (Figure 4). While a positive correlation also appears ($r=0.657$, p < 0.001), it is notably weaker than that observed for partial masking. This gap reinforces a key insight that full masking cannot exploit the fine-grained separability of activation distributions, whereas distribution-aware partial methods (e.g., APP, AURA) do.
5 Related Works
5.1 Sparse Autoencoders for Feature Discovery
Sparse Autoencoders (SAEs) have emerged as a powerful method for learning interpretable, monosemantic features from neural network activations Huben et al. (2024). Recent advances have focused on improving reconstruction quality and scaling through architectural and training strategy innovations such as JumpReLU activations, BatchTopK sparsity, gated, and end-to-end training frameworks Rajamanoharan et al. (2024a, b); Gao et al. (2025); Bussmann et al. (2024); Braun et al. (2024). Empirical analyses have validated SAEs’ ability to discover meaningful structures across different domains, from vision-language models to algorithmic patterns like temporal difference learning in LLMs Sun et al. (2025b); Pach et al. (2025); Demircan et al. (2025). Moreover, evaluation studies have highlighted both their utility for interpretability tasks and remaining challenges with polysemantic representations Kantamneni et al. (2025); Minegishi et al. (2025); Karvonen et al. (2025). However, none of these empirical analyses evaluated the separability of activation distributions in SAEs as a measure of polysemanticity.
5.2 SAE-Based Model Control
The interpretable features learned by SAEs enable precise control over language model behavior. Several works have demonstrated effective steering by carefully selecting and manipulating SAE features, with approaches ranging from supervised methods for identifying relevant dimensions to frameworks using hypernetworks Arad et al. (2025); He et al. (2025a, b); Bayat et al. (2025); Sun et al. (2025a); Minegishi et al. (2025). However, prior SAE-based control methods did not utilize posterior probabilities, limiting their precision.
5.3 Base Model Control and Causal Analysis
Complementing SAE-based approaches, researchers have developed techniques for direct activation control and causal analysis in the base language models. General frameworks for transporting activations facilitate intervention across model architectures Rodriguez et al. (2025), while causal tracing methods enable precise localization and editing of specific knowledge or biases Vig et al. (2020); Meng et al. (2022, 2023). These approaches offer foundational tools for probing and manipulating model behavior at the activation level. However, again, they do not leverage posterior probabilities for better intervention and have been applied exclusively to base models, not to representations learned by SAEs.
6 Conclusion
This work presents the first quantitative analysis of monosemanticity in SAEs compared to their dense base models through distributional lens. To better characterize monosemanticity, we introduce an activation distribution-aware concept separability score based on the Jensen–Shannon distance, which captures fine-grained distinctions in neuron activations across concepts. We also demonstrate that SAEs support more precise concept-level interventions than base models, particularly when using partial suppression. Building on this, we propose a new method, Attenuation via Posterior Probabilities, which achieves effective concept removal with least possible side effects.
7 limitations
To make density estimation computationally feasible at scale, APP replaces standard kernel density estimation (KDE) with a histogram-based approximation. While this approach substantially improves efficiency, it also introduces certain limitations. As the number of histogram bins increases, the accuracy of the estimated activation distributions improves, but so does the computational cost. Consequently, achieving the best possible performance of APP, in terms of precise density estimation and separability, requires significantly higher computational resources and runtime. Future work could explore alternative KDE methods to better balance accuracy and efficiency.
Appendix A Deepseek Concept Erasure Experiments
On DeepSeek, our method APP is highly competitive on the concept-removal metrics. Across all 20 comparisons of $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ (5 datasets $\times$ 2 metrics $\times$ Base/SAE), APP ranks as the best or second-best method in 13 cases. By contrast, the next strongest method, Range, achieves a top-two ranking in 12 out of 20 comparisons. This demonstrates that APP performs on par with or better than the existing approaches. Importantly, APP is also the least disruptive method, achieving the lowest increase in perplexity (DPPL) in 8 out of 10 DeepSeek experiments (all except POS for Base and POS for SAE). In one of these two exceptions, APP remains highly competitive with a perplexity increase of 0.269 vs. 0.253 for AURA. Overall, APP combines strong concept-removal performance with the smallest degradation in language modeling quality on DeepSeek.
Additionally, as shown in Table 3, applying interventions in the SAE representation yields larger $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ than in the Base representation across methods: for APP, $\Delta_{\mathrm{Conf}}$ increases on all five datasets and $\Delta_{\mathrm{Acc}}$ increases on 4 out of 5 (slightly lower on Emotions), indicating that SAEs promote greater concept separability and thereby enable more effective concept removal.
Appendix B NER Neurons Activation
As shown in Figure 5, the NER dataset exhibits patterns similar to AG News. Certain neurons in both the base model and SAEs (e.g., leftmost plots) show considerable overlap in their activation distributions across the four classes, indicating limited class discrimination. In contrast, neurons in the middle and right columns reveal more separable activation patterns.
Table 5: Concept Erasure Detailed metrics $D_{\mathrm{Acc}}$, $D^{\prime}_{\mathrm{Acc}}$, $D_{\mathrm{Conf}}$, and $D^{\prime}_{\mathrm{Conf}}$ for DeepSeek-R1.
Type Method POS AG News Emotions DBpedia NER
$D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$ $D_{\mathrm{Acc}}\uparrow$ $D^{\prime}_{\mathrm{Acc}}\downarrow$ $D_{\mathrm{Conf}}\uparrow$ $D^{\prime}_{\mathrm{Conf}}\downarrow$
Base APP 0.198 0.013 0.199 -0.011 0.193 0.009 0.363 -0.021 0.106 0.007 0.162 -0.029 0.160 0.005 0.242 -0.008 0.095 0.007 0.133 -0.010
Aura 0.081 0.032 0.103 0.034 0.029 0.014 0.125 0.054 0.099 0.032 0.179 0.033 0.241 0.030 0.348 0.067 0.051 0.013 0.074 0.003
Range 0.067 0.011 0.288 0.065 0.303 0.130 0.687 0.469 0.148 0.070 0.315 0.164 0.181 0.032 0.347 0.021 0.144 0.054 0.379 0.123
Adaptive 0.060 0.010 0.140 0.001 0.191 0.069 0.519 0.243 0.121 0.050 0.237 0.077 0.136 0.021 0.228 -0.016 0.121 0.031 0.236 0.039
Full 0.062 0.008 0.291 0.109 0.287 0.171 0.691 0.575 0.147 0.078 0.318 0.183 0.183 0.047 0.353 0.064 0.127 0.060 0.371 0.199
SAE APP 0.325 0.018 0.338 0.011 0.507 0.020 0.674 -0.134 0.098 0.006 0.339 -0.134 0.660 0.008 0.573 -0.085 0.581 0.014 0.838 -0.039
Aura 0.424 0.010 0.455 0.043 0.522 0.008 0.688 -0.086 0.098 0.009 0.339 -0.071 0.664 0.012 0.576 -0.198 0.604 0.015 0.869 -0.014
Range 0.420 0.067 0.479 0.057 0.497 0.044 0.644 -0.202 0.098 0.008 0.329 -0.329 0.515 0.017 0.347 -0.299 0.596 0.055 0.855 0.025
Adaptive 0.375 0.036 0.423 -0.008 0.465 0.032 0.607 -0.224 0.098 0.005 0.328 -0.324 0.499 0.011 0.337 -0.294 0.546 0.045 0.793 0.004
Full 0.422 0.177 0.481 0.174 0.501 0.301 0.652 0.174 0.098 0.008 0.329 -0.415 0.520 0.022 0.350 -0.326 0.603 0.303 0.867 0.378
Appendix C Details of Experiments
The detailed metrics ($D_{\mathrm{Acc}}$, $D^{\prime}_{\mathrm{Acc}}$, $D_{\mathrm{Conf}}$, and $D^{\prime}_{\mathrm{Conf}}$) before subtraction for Gemma-2-2b are reported in Table 4, and those for DeepSeek-R1 are presented in Table 5.
Appendix D Implementation Details
Hyperparameters. Range Masking, Adaptive Dampening, and Full Masking each rely on a saliency threshold hyperparameter that determines the top-$p\%$ of neurons considered most relevant to a given concept. For Gemma, we set $p=0.3$ for POS, AG News and NER, $p=0.2$ for Emotions and DBpedia. For DeepSeek, we use $p=0.3$, $0.4$, $0.1$, $0.2$, and $0.4$ for POS, AG News, Emotions, DBpedia, and NER, respectively. In contrast, AURA Suau et al. (2024) and our method, APP, do not require this hyperparameter. Both apply interventions across all neurons in the selected layer, avoiding the need to tune or justify a saliency cutoff. However, because SAE neurons typically activate on only a small subset of inputs for any given concept, we introduce an activation-frequency threshold $\tau$ to ensure reliability. Specifically, for each SAE neuron $h^{l}_{j}$, and each concept $c_{i}$, let $\mathcal{X}_{c_{i}}$ be the set of corresponding input samples. We define the firing frequency of $h^{l}_{j}$ on $c_{i}$ as
$f_{j,i}=\frac{\bigl|\{\,x\in\mathcal{X}_{c_{i}}:h^{l}_{j}(x)>0\}\bigr|}{|\mathcal{X}_{c_{i}}|}.$
We exclude neuron $h^{l}_{j}$ from all concept-erasure methods if $f_{j,i}<\tau$, as sparse activations preclude meaningful intervention. In all experiments, we set $\tau=0.1$.
Histogram-based KDE. To estimate the densities $f_{h^{l}_{j}\mid c_{i}}(x)$ required by APP, we use a kernel density estimation (KDE) implemented via a histogram-based approximation that preserves accuracy while greatly improving efficiency. During training, each neuron–concept activation distribution is discretized into $B$ uniform-width bins (we use $B=2048$) and we store the bin centers, counts, bandwidths, and normalization constants. At inference, a query activation $x$ is evaluated only against these $B$ centers rather than all $N$ training activations. Let $F$ be the number of neurons and $Q$ the number of query points; this reduces complexity from $\mathcal{O}(FNQ)$ time and $\mathcal{O}(FN)$ memory for naïve KDE to $\mathcal{O}(FBQ)$ and $\mathcal{O}(FB)$, respectively, yielding an approximate $N/B$-fold improvement in both inference speed and memory usage ($B\ll N$).
Intervention Location and Scope. All interventions are applied at a consistent computation point (immediately before the language-model head) to ensure comparability across models and methods. Specifically, interventions are introduced after the MLP and residual addition in the final transformer block. For SAEs, they are applied directly after the SAE activation nonlinearity. To isolate causal effects on prediction, interventions are restricted to the token corresponding to the model’s output label. This design allows precise measurement of each concept-erasure method’s influence on the final decision while avoiding confounding effects from earlier tokens.
Dataset Statistics. We report the dataset sizes and splits used in all experiments to ensure transparency and reproducibility. The AG News dataset contains 120,000 training and 7,600 test examples. The DBpedia dataset includes 100,000 training and 25,000 test examples, both randomly sampled using a fixed seed to ensure reproducibility. For Emotions, we use 16,000 training and 2,000 test examples. The POS Tagging and NER datasets each consist of 100,000 training and 10,000 test examples, also sampled with a fixed random seed.
Model Specifications. As described in the main text, we use the Gemma-2 model with 2 billion parameters (Hugging Face name: google/gemma-2-2b). For DeepSeek, we employ the DeepSeek-R1 model, specifically the distilled variant based on Llama-8B with 8 billion parameters (Hugging Face name: deepseek-ai/ DeepSeek-R1-Distill-Llama-8B).
Computation Details. All experiments were conducted using the university’s high-performance computing (HPC) cluster managed via the Slurm workload manager. We used NVIDIA A100 (40 GB) and V100 (32 GB) GPUs for activation collection and intervention process. Each job was allocated 1 GPU, 1–10 CPU cores, and 50 GB of RAM.
Result Reliability. All quantitative results reported in the concept-erasure tables correspond to a single representative run. To verify the stability of our findings, we reran selected datasets (e.g., AG News) multiple times and observed identical outcomes and consistent relative rankings across methods. For instance, the top-performing method in the initial run remained the best in repeated trials. These consistent results indicate that the reported values are reliable and reproducible.
Package Implementation and Parameter Settings. Our implementation leverages several standard Python and deep learning libraries. The Hugging Face Transformers package was used to load and run language models (AutoTokenizer, AutoModelForCausalLM). The transformer_lens and sae_lens libraries were employed for model inspection and sparse autoencoder integration. The scikit-learn library was used for evaluation metrics such as ROC-AUC, while pandas and numpy supported data processing. Default configurations were used for all libraries, with reproducibility ensured by fixing random.seed(42).
Figure 6: 65k SAEs show lower overlap than 16k, further supporting that greater capacity enables more distinct neuron-to-concept mappings.
Figure 7: JS distance vs. erasure ability for all partial erasure methods.
Appendix E All SAE Neurons Analysis
While our salient neuron analysis provides valuable insight into the most strongly responding neurons, it considers only a narrow slice of the activation space (specifically, the top 80 neurons per concept). This limited scope may miss neurons that, although not highly ranked by mean activation, are still consistently active across multiple concepts and contribute to polysemanticity. To address this limitation, we broaden our analysis to include all neurons that exhibit non-zero activation for any concept, offering a more comprehensive view of concept overlap beyond the most salient neurons. This extended analysis is conducted only for SAEs, since in the dense base model, all neurons are active across all inputs, rendering such overlap statistics uninformative. The results are shown in Figure 6, which reports the intersection-over-union of active neurons across concepts. By capturing both highly active and more subtly engaged neurons, this analysis reveals a complete picture of polysemantic behavior. Consistent with our earlier findings shown in Figure 1, we observe that higher-capacity SAEs (e.g., 65k dimensions) exhibit lower neuron overlap than their lower-capacity counterparts. Together, Figures 1 and 6 demonstrate that while SAEs significantly reduce polysemanticity, they do not eliminate it entirely. Polysemantic neurons remain present, albeit to a lesser extent.
Appendix F Concept Erasure and JS Distance
As illustrated in Figure 7, all partial concept erasure methods exhibit a strong Pearson correlation with our JS separability score metric. Specifically, each method achieves a correlation coefficient greater than 0.7 with statistically significant $p$-values ($p<0.01$), indicating a robust relationship between concept separability and erasure precision. These results suggest that partial erasure techniques can effectively leverage the separability inherent in activation representations to suppress the targeted concept while minimizing interference with unrelated ones. Among the four partial erasure approaches evaluated, AURA and APP demonstrate the highest correlations (0.765 and 0.744, respectively), highlighting their superior ability to exploit distributional distinctions between concept activations. We attribute this performance advantage to the fact that both AURA and APP explicitly model the distributions of both target and auxiliary concepts as discussed in Subsection 4.4 (Comparative Analysis of APP and AURA).
Appendix G JS Separability: DeepSeek vs. Gemma
Table 7 presents the JS separability scores computed for both the Gemma-2-2B and DeepSeek-R1 models across all datasets. As indicated in the table, every SAE variant consistently achieves higher separability scores than its respective base model. This consistent improvement confirms that the incorporation of SAEs enhances the distinction between concept activation distributions, leading to more interpretable internal representations. Moreover, when comparing SAEs with the same capacity, those with higher sparsity (corresponding to lower $l_{0}$ values) exhibit greater separability. These results confirms that sparsity plays a critical role in improving the distinctness of concept distributions.
ID Primary Class Related Class Sentence (excerpt) Base 1621 SAE-16k 332 SAE-65k 5997
1 Business Science/Tech Microsoft, Dell big winners of Air Force deal … for software services worth \$500M 12.18-12.18 $20.15$ $61.67$
2 Business Sports Retailers bet on poker’s rising popularity … playing cards and chips … holiday shopping season $0.52$ $26.31$ $45.32$
3 Business World Intellectual property rights … US companies in China … WTO commitments 4.27-4.27 $12.16$ $79.13$
4 Sports World Doping cases hit record … 24 athletes ousted for drug-related violations … Olympics … no US team member flunked a test $0.64$ $20.24$ $49.52$
5 Sports Science/Tech Suggs to practice … Cleveland Browns running back Lee Suggs cleared by doctors to return … still some haziness about his medical condition $5.14$ $26.99$ $54.39$
6 World Science/Tech Ivan skirts Grand Cayman, flooding homes … winds near 250 km/h … uprooting trees, bursting canals, and flooding homes as it churned toward Cuba 12.35-12.35 $19.02$ $62.22$
Table 6: Example AG News sentences illustrating cross-class activation patterns. Each sentence’s primary class is shown alongside a semantically related secondary class. Activation values reflect the response of three separable neurons identified in Figure 2.
POS AG News Emotions DBpedia NER
Gemma
Base 0.343 0.366 0.322 0.405 0.308
SAE
(width_16k/l0_116)
0.539 0.650 0.431 0.621 0.581
SAE
(width_16k/l0_285)
0.425 0.567 0.389 0.543 0.511
SAE
(width_65k/l0_93)
0.600 0.709 0.446 0.680 0.621
SAE
(width_65k/l0_197)
0.549 0.626 0.407 0.644 0.537
DeepSeek
Base 0.367 0.366 0.291 0.436 0.325
SAE 0.597 0.698 0.643 0.718 0.698
Table 7: JS separability score comparison across datasets for Gemma and DeepSeek. Bold values indicate the best score per dataset.
Appendix H Concept Distribution Analysis
We analyze concept-conditional activation distributions on the AG News dataset, focusing on three neurons with high class separability in Figure 2: Base Neuron 1621, SAE-16k Neuron 332, and SAE-65k Neuron 5997.
To examine how these neurons encode secondary topical cues, consider the following example from the AG News dataset:
“Microsoft, Dell big winners of Air Force deal … for software services … worth \$500 million …”
The primary label of this sentence is Business, though it clearly involves elements of Science/Tech. The per-neuron activations for this sentence are:
  • Base Neuron 1621: 12.1797-12.1797
  • SAE-16k Neuron 332: $20.1485$
  • SAE-65k Neuron 5997: $61.6663$
Across all three neurons, the activations fall within the Business class density in Fig. 2 but are shifted toward the Science/Tech density. This pattern is consistent with polysemantic encoding; the neurons primarily represent the Business topic (capturing the correct class) while remaining sensitive to Science/Tech signals (e.g., “Microsoft,” “software”), which pulls the activation toward the Science/Tech mode without leaving the Business distribution.
Table 6 extends the analysis to six mixed-topic sentences. In every case, the three neurons produce activations that sit within the primary class distribution yet are shifted toward the related class. Absolute scales differ across neurons, so values should be interpreted relatively against each neuron’s class-conditional densities in Fig. 2 rather than compared across neurons. Overall, the pattern supports a polysemantic account; these neurons track the primary topic while remaining sensitive to secondary topical signals.